# The Magnetic Ground State of Atacamite Cu2Cl(OH)3: The Crucial Role of Frustrated Zigzag Chains Revealed by Inelastic Neutron Scattering

### v1

@@ Abstract
I I. Introduction
Frustrated spin systems offer the distinct potential to host novel exotic magnetic phases [1, 2, 3, 4] and have become a platform through which to explore fundamental magnetism. The exotic phases emerge from specific magnetic interaction models. For instance, the corner-sharing tetrahedra of pyrochlore lattices can produce spin-ice states [5, 6, 3], and the kagomé lattice is expected to host a quantum spin-liquid state [2]. Identifying experimental systems that realize these models allows for rigorous testing of theoretical predictions under real-world conditions.
Herbertsmithite (ZnCu3(OH)6Cl2) [7] is one of the leading candidates for an experimental realization of a quantum spin-liquid material with its geometrically perfect spin- kagomé lattice. It sits within the broader atacamite family of minerals [8], with the general formula (ZnxCu4-x(OH)6Cl2) [9] encompassing polymorphs with various ratios of Zn-Cu substitution.
Atacamite (Cu2Cl(OH)3) refers to the orthorhombic polymorph of this mineral family, with full Cu2+ occupation. The spin- Cu2+ sites class atacamite as a quantum-magnetic system, and experiments have characterized its geometric frustration effects [10]. Furthermore, atacamite exhibits a complex magnetic phase diagram, including an antiferromagnetic ground state at = 8.9 K [11] (with a magnetic propagation vector of = [, 0, ] [12] and with suppressed magnetic momeents of and for Cu(1) and Cu(2) sites, respectively.) and a sequence of metamagnetic transitions up to 30 T, including a plateau-like behvaior of the bulk magnetization above 31.5 T (H axis) [11]. However, the precise nature of these transitions and phases remained ambiguous.
@@
Regarding the magnetic interactions driving the ground-state magnetism of atacmite, recent density functional theory (DFT) calculations indicate that the dominant exchange interactions form a weak 3D network of antiferromagnetic sawtooth chains [11]. The five largest exchange pathways are shown in Fig. 1, with the sawtooth chain running along the direction. Interest has followed atacamite not only for its compositional similarities to Herbertsmithite, but also due to the unmatched opportunity it offers to experimentally investigate the complexities of a low-dimensional quantum spin system within the intensely-studied sawtooth chain model [14, 15, 16, 17, 18, 19, 20, 21].
Although the sawtooth chain is the dominant coupling motif in atacamite, DFT calculations also reveal a weaker perpendicular chain formed by competing nearest and next-nearest interactions [11]. This is the S= zigzag spin chain [22, 23, 24, 25] (hereafter referred to as simply the ‘zigzag chain’ to avoid confusion with our exchange-label naming convention), with Hamiltonian
@@
In this investigation, we demonstrate that atacamite’s zigzag chain plays a previously overlooked role in its magnetic behavior at low temperatures. We present the first investigation of atacamite’s spin-wave excitation spectrum using inelastic neutron scattering (INS) within its ordered phase to directly probe its magnetic interactions. We find that the zigzag chain is crucial for modeling atacamite, and, when combined with a mean-field treatment, an effective-model Hamiltonian for atacamite is produced that agrees with INS measurements in the low-energy regime. Furthermore, this agreement stands in contrast to the poor match obtained using the full DFT-derived model, whose dominant sawtooth chain is of a much higher energy scale than the zigzag chain. An effective separation of energy scales between the two chains is thus presented as a puzzling characteristic of a material realization of a frustrated quantum magnet.
The compatibility between a magnetic model and INS measurements can only be justified within a robust theoretical framework. Linear spin-wave theory (LSWT) stands out as a straightforward, common [30], and reliable method for making quantitative comparisons between a magnetic Hamiltonian and INS experiments in long-range ordered spin systems. The surprising success of LSWT in its application to quantum systems is rooted in the abundance of semiclassical behaviors [31], even in the limit [32]. The LSWT calculations within this work were performed using the Sunny package [33] where the magnetic structure refined from neutron diffraction data [11] was chosen for the spin structure input.
II II. Experimental Methods
LSWT and INS measurements are often used together to validate first-principles calculations and refine empirical magnetic models, including for frustrated (quantum) spin systems [5, 34, 35, 36, 37, 38]. Single-crystal neutron scattering measurements of atacamite’s spin-wave spectrum were performed at the Australian Nuclear Science and Technology Organisation (ANSTO) on the Pelican [39] time-of-flight INS spectrometer with the sample rotated in the H0L plane. Pelican was configured with 4.69 Å neutrons and an approximate resolution of 0.135 meV at the elastic line.
@@
Data was collected for the sample at 1.5 K, 6 K, and 20 K in a closed-cycle helium cryostat, and was measured across a sample rotation of in intervals. A vanadium sample was used to normalize the detector efficiencies and an empty cryostat scan was used for background subtraction.
III III. Results
@@
The INS data along the H-reciprocal-lattice direction, shown in Fig. 2(a), is a 2D cut where neutron counts have been integrated within the interval [–2.5,2.5] along L, and [–0.2,0.2] along K. The data within the volume along L is relatively non-dispersive, and the intensity is distributed over a broadened energy range, as shown in Section VI of the supplemental material [41]. The measurements probe the low-energy region of atacamite’s spin-wave excitations with an upper-energy threshold of 2.5 meV.
@@
Along the H direction, two distinct, dispersive, energy-offset spin-wave modes are observed. A linear energy cut at H = 0.5 r.l.u. is shown in Fig. 3 and is taken with the same integration widths in L and K as the H-dependent spectra in Fig. 2(a), with an integration width of 0.04 r.l.u. in H. These data reveal two strong peaks in intensity – one at 1.7 meV and one at 1.3 meV. The lower-energy 1.3 meV peak is broad, and the neutron count doesn’t reach the background level at the magnetic zone center until approximately 0.75 meV, as shown in Fig. 4 (see Fig. S10 [41] for more details on the gap quantification). Atacamite thus has an energy gap of at least 0.75 meV along the H00 direction, but, due to instrumental broadening, could be as high as the peak center of meV, which would be compatible with the gap extracted from heat capacity measurements [13].
@@
LSWT calculations performed using the five largest exchange interactions yielded by previous DFT results [11], which includes both the sawtooth and zigzag chains, fail to accurately reproduce all characteristics of these INS measurements. Specifically, spin-wave excitations appear at higher energies than are observed in the INS measurements, the asymmetry of the mode intensities do not match, and the lower-energy mode is not gapped (for further LSWT details for the full-DFT coupling network, see Figs. S5 and S6 [41]).
@@
These shortcomings demonstrate that the full-DFT model does not adequately capture the magnetism governing atacamite’s low-energy excitations within a semiclassical approach. The discrepancies in the full-DFT model may stem from a the simplified treatment of quantum effects in LSWT, and could be alleviated in a fully quantum-mechanical treatment. Unfortunately, frustration and the three-dimensional nature of the full model make such treatment unfeasible. As an alternative, for a given energy scale, we can construct an effective model, which treats a subset of relevant exchanges exactly, while other exchanges enter the model in a form of an effective field [22, 23, 24, 25]. In atacamite, we achieve this by a mean-field treatment of the antiferromagnetic zigzag chain model along the axis using the Hamiltonian
@@
The parameters of this model have been refined through least-squares fitting of LSWT calculations to the INS data. The refinement arrives at = 0.90 meV, = 1.91 meV, and = 0.255 meV, giving a characteristic / ratio of 2.12. Classical LSWT calculations of the antiferromagnetic zigzag chain with ratios above 1 also match exact diagonalization calculations (see Fig. S11 in the supplemental information [41]), suggesting quantum effects are negligible for these calculations.
The refined LSWT scattering function calculation is shown in Fig. 2(b). The result shows a remarkable qualitative match with experimental data along the H direction, namely the spin-wave amplitudes, the relative intensities of the dispersion, asymmetry in the dispersion intensities, the weakening intensity towards the magnetic zone boundary, the spin-wave frequencies, and, crucially, the spin-wave gap. A 1D-cut of the calculated spin-wave spectra further confirms an excellent match with the experimental peak positions and gap size, as shown in Fig. 3.
@@
Without frustration, a zigzag chain is expected to adopt an incommensurate spiral order which is nearly of four-fold periodic [22]. The fact that the ground state is exactly four-fold commensurate implies that weak quantum fluctuations have selected this spin state through order-by-disorder from a frustration-induced classically degenerate set. This breaks symmetry, explaining the large observed gap [45].
@@
IV IV. Discussion
Our results have revealed the critical nature of in modeling atacamite, and INS has allowed us to refine its highly -sensitive value. These results, however, do not directly cast doubt on the validity of the previous DFT approach; the previous calculations correspond to the energy scale of the Curie-Weiss temperature [11], their dominant couplings agree with the Goodenough-Kanamori-Anderson rule [48, 49, 50] by association with atacamite’s largest Cu-O-Cu superexchange pathways, and they contain the very zigzag chain geometry used for our model.
@@
Secondly, the energy scales between the zigzag and sawtooth chain are effectively separated, reflecting different roles in atacamite’s spin dynamics. Although all isotropic couplings can be deduced simultaneously through DFT, low-energy spin-wave calculations performed with the full coupling scheme are unable to fully reproduce INS observations. Meanwhile, at temperatures far above TN = 8.9 K, the Curie-Weiss temperature derived from inverse susceptibility aligns with the dominant energy scale of the sawtooth chain. Additionally, magnetic order is suppressed near the energy scale of the zigzag chain [13], suggesting it governs atacamite’s ordering dynamics.
@@
In contrast, while prior DFT calculations for the dominant direct exchange couplings are consistent with high-temperature experimental results, they fail to account for the low-energy spin dynamics below the ordering temperature in a semiclassical LSWT treatment. This suggests that the full isotropic-exchange DFT model may be missing the subtleties required to capture the low-energy magnetic dynamics in a frustrated magnet. Frustration also enhances quantum effects, suppressing weaker interactions such as interchain couplings. This has led to an effective energy scale separation between two distinct 1D-chain models. Nevertheless, by employing an effective model of a field-moderated zigzag chain, we successfully capture the essential features of the experimental dispersion relations and provide a reliable description of atacamite’s low-energy magnetism. This way, we have uncovered the frustrated interchain exchange in atacamite leading to long-range magnetic order below its Néel temperature, and where the magnetic state is selected through order-by-disorder. In the light of the recent experimental results in high magnetic fields [13], it seems to be this frustrated interchain coupling network which is effectively broken up when the Cu(2) spins are fully field polarized in magnetic fields beyond the quantum critical point. Numerical results here supported the scenario of an (effective) dimensional reduction of the spin system upon which long-range magnetic order is suppressed.
@@
VI Acknowledgments
This research used the Pelican time-of-flight spectrometer for proposal number P8226 at ACNS, a part of ANSTO, the Australian Nuclear Science and Technology Organisation. J. L. Allen would like to thank AINSE Limited for providing financial assistance (Award - PGRA) to enable this research. The authors acknowledge and thank A. U. B. Wolter from The Leibniz Institute for Solid State and Materials Research for her extensive contributions to atacamite research, and for constructive discussions while preparing the manuscript.

### v2

@@ Abstract
I Introduction
Frustrated spin systems offer the distinct potential to host novel exotic magnetic phases [1, 2, 3, 4] and have become a platform through which to explore fundamental magnetism. The exotic phases emerge from specific magnetic interaction models. For instance, the corner-sharing tetrahedra of pyrochlore lattices can produce spin-ice states [5, 6, 3], and the kagome lattice is expected to host a quantum spin-liquid state [2]. Identifying experimental systems that realize these models allows for rigorous testing of theoretical predictions under real-world conditions.
Herbertsmithite (ZnCu3(OH)6Cl2) [7] is one of the leading candidates for an experimental realization of a quantum spin-liquid material with its geometrically perfect spin- kagome lattice. It sits within the broader atacamite family of minerals [8], with the general formula ZnxCu4-x(OH)6Cl2 [9] encompassing polymorphs with various ratios of Zn-Cu substitution.
Atacamite (Cu2Cl(OH)3) refers to the orthorhombic polymorph of this mineral family, with full Cu2+ occupation. The spin- Cu2+ sites class atacamite as a quantum-magnetic system, and experiments have characterized its geometric frustration effects [10]. Furthermore, atacamite exhibits a complex magnetic phase diagram, including an antiferromagnetic ground state at = 8.9 K [11] (with a magnetic propagation vector of = [, 0, ] [12] and with suppressed magnetic moments of and for Cu(1) and Cu(2) sites, respectively) and a sequence of metamagnetic transitions up to 30 T, including a plateau-like behavior of the bulk magnetization above 31.5 T (H axis) [11]. However, the precise nature of these transitions and phases remained ambiguous.
@@ I Introduction
Regarding the magnetic interactions driving the ground-state magnetism of atacmite, recent density functional theory (DFT) calculations indicate that the dominant exchange interactions form a weak 3D network of antiferromagnetic sawtooth chains [11]. The five strongest exchange pathways ( meV, meV, meV, meV, and meV) are shown in Fig. 1, with the sawtooth chain running along the direction. Interest has followed atacamite not only for its compositional similarities to Herbertsmithite, but also due to the unmatched opportunity it offers to experimentally investigate the complexities of a low-dimensional quantum spin system within the intensely-studied sawtooth chain model [14, 15, 16, 17, 18, 19, 20, 21].
Although the sawtooth chain is the dominant coupling motif in atacamite, DFT calculations also reveal a weaker perpendicular chain formed by competing nearest and next-nearest interactions [11]. This is the S= zigzag spin chain [22, 23, 24, 25] (typically referred to as the zigzag chain in the literature, and hereafter referred to as simply the ‘zigzag chain’ to avoid confusion with our exchange-label naming convention), with Hamiltonian
@@
In this investigation, we demonstrate that atacamite’s zigzag chain plays a previously overlooked role in its magnetic behavior at low temperatures. We present the first investigation of atacamite’s spin-wave excitation spectrum using inelastic neutron scattering (INS) within its ordered phase to directly probe its magnetic interactions. Similar approaches have been successfully employed to constrain and determine complex Hamiltonians in other zigzag-chain materials [30, 31]. We find that the zigzag chain is crucial for modeling atacamite, and, when combined with a mean-field treatment, an effective-model Hamiltonian for atacamite is produced that agrees with INS measurements in the low-energy regime. Furthermore, this agreement stands in contrast to the poor match obtained using the full DFT-derived model, whose dominant sawtooth chain is of a much higher energy scale than the zigzag chain. An effective separation of energy scales between the two chains is thus presented as an interesting characteristic of a material realization of a frustrated quantum magnet.
The compatibility between a magnetic model and INS measurements can only be justified within a robust theoretical framework. Linear spin-wave theory (LSWT) stands out as a straightforward, common [32], and reliable method for making quantitative comparisons between a magnetic Hamiltonian and INS experiments in long-range ordered spin systems.
LSWT and INS measurements are often used together to validate first-principles calculations and refine empirical magnetic models, including for frustrated (quantum) spin systems [5, 33, 34, 35, 36, 37].
The surprising success of LSWT in its application to quantum systems is rooted in the abundance of semiclassical behaviors [38], even in the limit [39] [30]. The LSWT calculations within this work were performed using the Sunny package [40] where the magnetic structure refined from neutron diffraction data [11] was chosen for the spin structure input.
II Experimental Methods
Single-crystal neutron scattering measurements of atacamite’s spin-wave spectrum were performed at the Australian Nuclear Science and Technology Organisation (ANSTO) on the Pelican [41] time-of-flight INS spectrometer with the sample rotated in the H0L plane. Pelican was configured with 4.69 Å neutrons and an approximate resolution of 0.135 meV at the elastic line.
@@ II Experimental Methods
Data were collected for the sample at 1.5 K, 6 K, and 20 K in a closed-cycle helium cryostat, and were measured across a sample rotation of in intervals. A vanadium sample was used to normalize the detector efficiencies and an empty cryostat scan was used for background subtraction.
III Results
@@
The INS data along the H-reciprocal-lattice direction, shown in Fig. 2a, is a 2D cut where neutron counts have been integrated within the interval [–2.5,2.5] along L, and [–0.2,0.2] along K. The data within the volume along L is relatively non-dispersive, and the intensity is distributed over a broadened energy range, as shown in Section VI of the supplemental material [43]. The measurements probe the low-energy region of atacamite’s spin-wave excitations with an upper-energy threshold of 2.5 meV.
@@
Along (H,0,0), two distinct, dispersive, energy-offset spin-wave modes are observed. A linear energy cut at H = 0.5 r.l.u. is shown in Fig. 3 and is taken with the same integration widths in L and K as the H-dependent spectra in Fig. 2a, with an integration width of 0.04 r.l.u. in H. These data reveal two strong peaks in intensity – one at 1.7 meV and one at 1.3 meV. The lower-energy 1.3 meV peak is broad, and the neutron count doesn’t reach the background level at the magnetic zone center until approximately 0.75 meV, as shown in Fig. 4 (see Fig. S10 [43] for more details on the gap quantification). Atacamite thus has an energy gap of at least 0.75 meV along the H00 direction, but, due to instrumental broadening, could be as high as the peak center of meV, which would be compatible with the gap extracted from heat capacity measurements of approximately 1 meV [13].
@@
LSWT calculations performed using the five strongest exchange interactions yielded by previous DFT results [11], which includes both the sawtooth and zigzag chains, fail to accurately reproduce all characteristics of these INS measurements. Specifically, spin-wave excitations appear at higher energies than are observed in the INS measurements, the asymmetry of the mode intensities do not match, and the lower-energy mode is not gapped (for further LSWT details for the full-DFT coupling network, see Figs. S5 and S6 [43]).
@@
These shortcomings demonstrate that the full-DFT model does not adequately capture the magnetism governing atacamite’s low-energy excitations within a semiclassical approach. The discrepancies in the full-DFT model may stem from the simplified treatment of quantum effects in LSWT, and could be alleviated in a fully quantum-mechanical treatment. Unfortunately, frustration and the three-dimensional nature of the full model make such treatment unfeasible. As an alternative, for a given energy scale, we can construct an effective model, which treats a subset of relevant exchanges exactly, while other exchanges enter the model in a form of an effective field [22, 23, 24, 25]. In atacamite, we achieve this by a mean-field treatment of the antiferromagnetic zigzag chain model along the axis using the Hamiltonian
@@
The parameters of this model have been refined through a particle swarm algorithm comparing LSWT calculations to the INS data to find the global minimum. The refinement arrives at = 0.90 meV, = 1.91 meV, and = 0.255 meV, giving a characteristic / ratio of 2.12. Classical LSWT calculations of the antiferromagnetic zigzag chain with ratios above 1 are qualitatively similar to exact diagonalization calculations (see Fig. S11 in the supplemental information [43]), suggesting quantum effects do not preclude semiclassical calculations from adequately characterizing the excitation spectrum in this system.
The refined LSWT scattering function calculation is shown in Fig. 2b. The result shows a remarkable qualitative match with experimental data along (H,0,0), namely the spin-wave amplitudes, the relative intensities of the dispersion, asymmetry in the dispersion intensities, the weakening intensity towards the magnetic zone boundary, the spin-wave frequencies, and, crucially, the spin-wave gap. A constant- cut of the calculated spin-wave spectra further confirms an excellent match with the experimental peak positions and gap size, as shown in Fig. 3.
@@
Without frustration, a zigzag chain is expected to adopt an incommensurate helical order which is nearly four-fold periodic [22]. The fact that the ground state is exactly four-fold commensurate implies that weak quantum fluctuations have selected this spin state through order-by-disorder from a frustration-induced classically degenerate set. This breaks symmetry, explaining the large observed gap [47].
@@
IV Discussion
Our results have revealed the critical nature of in modeling atacamite, and INS has allowed us to refine its highly -sensitive value. These results, however, do not directly cast doubt on the validity of the previous DFT approach; the previous calculations correspond to the energy scale of the Curie-Weiss temperature [11], their dominant couplings agree with the Goodenough-Kanamori-Anderson rule [50, 51, 52] by association with atacamite’s strongest Cu-O-Cu superexchange pathways, and they contain the very zigzag chain geometry used for our model.
@@ IV Discussion
Secondly, the energy scales between the zigzag and sawtooth chain are effectively separated by one order of magnitude, reflecting different roles in atacamite’s spin dynamics. Although all isotropic couplings can be deduced simultaneously through DFT, low-energy spin-wave calculations performed with the full coupling scheme are unable to fully reproduce INS observations. Meanwhile, at temperatures far above T = 8.9 K, the Curie-Weiss temperature of 130 K derived from inverse susceptibility [12] aligns with the dominant energy scale of the sawtooth chain [11]. Additionally, magnetic order is suppressed near the energy scale of the zigzag chain [13], suggesting it governs atacamite’s ordering dynamics.
@@ V Conclusions
In contrast, while prior DFT calculations for the dominant direct exchange couplings are consistent with high-temperature experimental results, they fail to account for the low-energy spin dynamics below the ordering temperature in a semiclassical LSWT treatment. This suggests that the DFT+ approach is unable to describe the magnetism across different energy scales with a single value in this complex and frustrated magnetic system. Frustration also enhances quantum effects, suppressing weaker interactions such as interchain couplings. This has led to an effective energy scale separation between two distinct 1D-chain models. Nevertheless, by employing an effective model of a field-moderated zigzag chain, we successfully capture the essential features of the experimental dispersion relations and provide a reliable description of atacamite’s low-energy magnetism. This way, we have uncovered the frustrated interchain exchange in atacamite leading to long-range magnetic order below its Néel temperature, and where the magnetic state is selected through order-by-disorder. In the light of the recent experimental results in high magnetic fields [13], it seems to be this frustrated interchain coupling network which is effectively broken up when the Cu(2) spins are fully field polarized in magnetic fields beyond the quantum critical point. Numerical results here supported the scenario of an (effective) dimensional reduction of the spin system upon which long-range magnetic order is suppressed.
@@
Acknowledgments
This research used the Pelican time-of-flight spectrometer for proposal number P8226 at ACNS, a part of ANSTO, the Australian Nuclear Science and Technology Organisation. J. L. Allen would like to thank AINSE Limited for providing financial assistance (Award - PGRA) to enable this research. O.J. and S.N. were supported by the German Forschungsgemeinschaft (DFG, German Research Foundation) through SFB 1143 (Project ID 247310070). The authors acknowledge and thank A. U. B. Wolter from The Leibniz Institute for Solid State and Materials Research for her extensive contributions to atacamite research, and for constructive discussions while preparing the manuscript. Finally, we thank Ulrike Nitzsche for technical assistance.